# Towards Explainable Bilingual Multimodal Misinformation Detection and Localization

### v1

@@ 4 BiMi: A Bilingual Multimodal Misinformation Detection Framework
Figure 4: The overview of the training strategy. Three stages: domain alignment on news data, instruction tuning with task-specific prompts, and GRPO optimization with structured rewards. Each stage builds upon the previous one, progressively enhancing task performance.
@@ 4.1 Framework
Modeling. We adopt Gemma 3 [39] as the pre-trained model of BiMi due to its strong multilingual understanding and cross-modal reasoning capabilities. Gemma 3 is a multilingual MLLM with strong vision-language alignment and cross-lingual reasoning, making it well-suited for detecting subtle manipulations across images and bilingual subtitles. In our framework, the input image is overlaid with Chinese and English subtitles to form a unified visual input, which is then processed by a vision encoder to extract patch-level embeddings. These representations are fused through learned attention layers to form a unified multimodal context. This design allows the model to jointly reason over visual and bilingual textual cues, making it well-suited for detecting subtle manipulations such as object removal, semantic shifts, and cross-lingual inconsistencies.
Retrieval Module. To improve adaptability to emerging misinformation, BiMi integrates a retrieval module at inference time. We first extract Chinese and English subtitles from the image using OCR, then query the Google Search API [42] and collect the top-3 results. Retrieved texts are appended to the input prompt, providing up-to-date context that helps the model assess semantic consistency. This retrieval-augmented setup enables BiMi to generalize better to previously unseen or rapidly evolving misinformation cases.
@@ 4.2 Multi-stage Training
Stage 1: Domain Alignment. To adapt the model to the linguistic and visual characteristics of the news domain, we perform instruction-based tuning using image-caption pairs from VisualNews [41]. Each instance consists of an instruction prompt (e.g., Describe this image in English/Chinese) and a corresponding caption in either English or Chinese. We use captions from the Washington Post subset and translate each English caption into Chinese to ensure cross-lingual coverage. The model learns to generate captions conditioned on both the image and the prompt, thereby establishing strong visual-textual grounding across both languages. This stage enhances the model’s ability to handle real-world journalistic content and supports bilingual reasoning in downstream misinformation detection tasks.
@@
[…]
@@
[…]
@@
where $x_{i}$ denotes the $i$-th token in the assistant response. Only the tokens in $X_{\text{a}}$ are used to compute the loss. This tuning step encourages the model to align with the structure and reasoning required for fine-grained misinformation detection.
Stage 3: GRPO-based Reasoning Optimization. To improve the model’s reasoning quality and explanation consistency, we apply Group Relative Policy Optimization (GRPO) [12], a reinforcement learning strategy that adjusts the policy by comparing outputs within a batch rather than relying on absolute rewards alone.
@@
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}\left[{\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta_{\text{old}}}(q)}\right]\frac{1}{N}\sum_{i=1}^{N}\left\{\min\left[s_{1}\cdot A_{i},\;s_{2}\cdot A_{i}\right]-\beta\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right\}$ (2)
$A_{i}$ $=\frac{r_{i}-\mathrm{mean}\{r_{1},r_{2},\ldots,r_{N}\}}{\mathrm{std}\{r_{1},r_{2},\ldots,r_{N}\}},\quad s_{1}=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\quad s_{2}=\mathrm{clip}\left(s_{1},\;1+\epsilon,\;1-\epsilon\right),$ (3)
@@
Format reward. To enforce structured outputs, we define a format reward $R_{\text{format}}$ that checks whether the model response follows a predefined format, using <answer></answer> tags for predictions and an optional <think></think> section for explanations. This ensures consistency across outputs and simplifies automatic evaluation. We use regular expressions for validation: $R_{\text{format}}=\begin{cases}0,&\text{if output matches expected format}\\ 1,&\text{otherwise}\end{cases}$.
Localization reward. We use an IoU-based reward $R_{\text{Loc}}$ to assess the alignment between predicted and ground-truth tampered regions. To simplify computation, both are represented as bounding boxes. The reward is defined as: $R_{\text{Loc}}=\frac{|\mathcal{M}_{\text{pred}}\cap\mathcal{M}_{\text{gt}}|}{|\mathcal{M}_{\text{pred}}\cup\mathcal{M}_{\text{gt}}|}$, where $\mathcal{M}_{\text{pred}}$ and $\mathcal{M}_{\text{gt}}$ denote the predicted and ground-truth bounding boxes, respectively.
Classification reward. To encourage accurate decision-making in multimodal misinformation detection, we define a classification reward $R_{\text{cls}}$ that compares the model’s predicted label with the ground-truth annotation. This applies to tasks such as detecting whether the image is manipulated or whether the subtitles are semantically consistent. $R_{\text{cls}}=\begin{cases}1,&\text{if }C_{\text{pred}}=C_{\text{gt}}\\ 0,&\text{otherwise}\end{cases}$.
The final reward combines all task-specific objectives, including formatting, classification, localization: $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$. This unified reward encourages the model to generate structured, accurate, and interpretable predictions across modalities.
@@ 5.1 Experimental Setup
We evaluate BiMi on BiMiBench and MMFakeBench [8]. On BiMiBench, the model performs six-class classification, tampered region localization, and explanation generation. For MMFakeBench, which lacks subtitle information, we follow its four-class setting and provide text context via prompts.
Implementation Details. We use the Gemma 3 as the LLM, applying our three-stage post-training pipeline. All experiments are run on 48GB V100 GPUs with a batch size of 16, learning rate $1\times 10^{-5}$, and frozen vision encoder. GRPO is applied for two additional epochs. We design structured instruction prompts to guide the model in tasks, as detailed in Table 3.
Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags:
<think>Your explanation here.</think>
<answer>{"classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}]}</answer>
@@
Baselines. We compare BiMi against a range of strong multimodal baselines, including InternVL3 [27], Qwen3 [47], LLaVA-1.6 [33], and LLama3.1 [48], which represent leading approaches in vision-language modeling. We also include specialized misinformation detection systems such as SNIFFER [7] and MMD-Agent [8], designed for out-of-context and multimodal misinformation. All models are evaluated using the same inputs: the image with overlaid Chinese-English subtitles and a task-specific prompt. For ablation, we test BiMi variants without GRPO, without retrieval, and with alternative MLLM backbones.
Evaluation Metrics. We report Accuracy (ACC) and F1 score for classification, IoU for localization. Inspired by [30], we evaluate explanation quality via BERTScore [49] between response and reference explanations generated by prompting GPT-4o [50] with labels and manually verified for correctness.
Table 4: Main results on BiMiBench. BiMi achieves the best performance across all metrics. Bold and underline denote best and second-best results.
Method LLM Res. ACC F1 IoU BERT
InternVL3 [27] Qwen2.5-7B 224 20.85 18.42 7.23 71.82
InternVL3 [27] Qwen2.5-14B 224 28.47 26.71 14.38 76.54
Qwen3 [47] Qwen3-8B 224 21.43 14.96 6.28 73.92
LLaVA-1.6 [33] Vicuna-7B 224 22.91 20.74 8.91 75.77
LLama3.1 [48] Llama-3.1-8B 224 24.61 21.94 5.27 72.91
Gemma3 [48] Gemma3-4B-IT 336 17.48 11.73 4.29 70.84
SNIFFER [7] Vicuna-13B 224 37.90 33.63 80.41
MMD-Agent [8] Qwen2.5-7B 224 34.78 29.48
BiMi Gemma3-4B-IT 336 46.80 42.79 30.24 82.90
@@ 5.2 Main Results
Table 5: MMFB results.
Method ACC F1
InternVL3 42.69 34.80
Qwen3 44.82 39.49
LLaVA-1.6 30.09 24.31
LLama3.1 32.67 28.30
Gemma3 27.19 22.04
SNIFFER 67.49 61.44
MMD-Agent 62.78 52.83
BiMi 70.94 62.49
Performance Comparison. We compare BiMi with strong baselines across three subtasks on the BiMiBench test set: (1) misinformation detection, (2) tampered region localization, and (3) explanation generation. Results are summarized in Table 4. BiMi significantly outperforms all baselines in both all tasks. In particular, it achieves a +8.9 accuracy gain in image-subtitle consistency, a +15.9 localization accuracy, and a +2.5 BERTScore improvement in explanation generation over the strongest baseline. BiMi’s performance also benefits from the GRPO reward design, which improves explanation specificity and localization precision. The results on MMFakeBench (4-class classification task) are shown in Table 5.
Ablation Study. We investigate the contributions of each stage in our pipeline using an ablation study. (i) Domain Alignment. Removing this stage reduces accuracy by 2.35 and BERTScore by 4.48, indicating its role in grounding visual-textual reasoning on news content. (ii) Instruction Tuning. Eliminating instruction tuning leads to a drastic performance drop, accuracy falls by 31.18 and BERTScore by 12.06, highlighting its importance in enabling structured bilingual understanding. (iii) GRPO. Without GRPO, the model fails to generate specific explanations tied to image or
@@
subtitle content, reducing accuracy by 5.7 and BERTScore by 8.09. (iv) Output Order. Reversing the output order (classify before explain) slightly lowers accuracy (–1.94) and BERTScore (–1.11), suggesting that generating explanations first facilitates more coherent reasoning. Table 6 summarizes the results. These findings confirm that all components are essential, with instruction tuning and GRPO being particularly critical for boosting accuracy and explanation quality.
@@ 5.3 Explainability Analysis
Beyond accurate classification and localization, generating high-quality natural language explanations is crucial for interpretable misinformation detection. We analyze explanation ability through (1) qualitative comparisons, (2) the influence of the retrieval module, and (3) evaluation under bilingual.
Explanation Quality. To understand BiMi’s reasoning ability, we compare its output with InternVL3 [27] on a representative case (Figure 5). Both models recognize that the subtitles are semantically aligned, showing bilingual understanding. However, InternVL3 relies on external factual cues and overlooks the mismatch between image and text. In contrast, BiMi grounds its decision in the image, correctly identifies the manipulation, and localizes the tampered region. It also provides a clearer, step-by-step explanation, showing stronger visual grounding and more faithful reasoning.
Figure 5: Comparison of explanations from InternVL3 (middle) and our model (right). The top left is input. The bottom left is the original sample. Note: Due to space constraints, some model responses are truncated. Full examples and additional results are provided in the Appendix.
Effect of Retrieval. We analyze a real-world social media post with subtle bilingual inconsistency (Figure 6). The removed "student" in the Chinese translation alters the meaning. Without retrieval, the model misclassifies due to a lack of context. The retrieval result provides a key clue for classification. BiMi correctly identifies the inconsistency, showing its benefit for nuanced cross-lingual reasoning.
@@
Bilingual multimodal capability. To evaluate bilingual understanding, we construct a variant of the test set where the model receives only one subtitle, Chinese or English. This reduces the task to a 4-category classification. By ignoring cross-lingual consistency, the single-language setting reduces reasoning difficulty. As shown in Table 7, the model performs better with English-only inputs, likely due to the predominance of English data during training. This highlights a gap in bilingual support and the need for stronger Chinese-language modeling in this domain.
Figure 6: Social media example with subtle bilingual inconsistency. The Chinese translation omits "student", changing the meaning and potentially misleading Chinese readers.
Failure Cases. We observe that BiMi occasionally fails when manipulations are extremely subtle or when OCR fails to extract accurate subtitles. In such cases, either the localization is off-target, or the explanation is overly general. More failure cases and analysis are provided in the Appendix.

### v3

@@ 4 BiMi: A Bilingual Multimodal Misinformation Detection Framework
Figure 4: The overview of the training strategy. Three stages: domain alignment on news data, instruction tuning with task-specific prompts, and GRPO optimization with structured rewards.
@@ 4.1 Framework
Modeling. We adopt Gemma 3 [team2025gemma] as BiMi’s backbone for its strong multilingual understanding and vision–language alignment, enabling detection of subtle manipulations across images and bilingual subtitles. The input image with overlaid Chinese and English subtitles is encoded into patch-level embeddings, which are fused through attention layers to create a unified multimodal context. This design allows joint reasoning over visual and bilingual textual cues to identify object manipulation, semantic shifts, and cross-lingual inconsistencies.
Retrieval Module. To enhance adaptability to emerging misinformation, BiMi employs a retrieval module at inference. Chinese and English subtitles are extracted via OCR to form a bilingual query to the Google Search API [googleapi2024]; the top-3 retrieved passages are prepended to the input as a unified prompt $\mathcal{P}=\operatorname{concat}(R,I,S)$, where $R$, II, and $S$ denote the retrieved context, image, and subtitles. This auxiliary context supplies timely external information that improves the model’s generalization to previously unseen or rapidly evolving misinformation, while the final predictions remain grounded in the image–subtitle content.
@@ 4.2 Multi-stage Training
Stage 1: Domain Alignment. To adapt the model to the linguistic and visual traits of the news domain, we perform instruction-based tuning on VisualNews [liu-etal-2021-visual] image–caption pairs. Each instance uses an instruction prompt (e.g., Describe this image in English/Chinese) with an English or translated Chinese caption. This trains the model to generate captions from images and prompts, strengthening visual–text grounding and bilingual reasoning for downstream misinformation detection.
@@
$X=\texttt{<system>}\,[X_{\text{q}};X_{v}]\,\texttt{<STOP>}\,X_{\text{a}}\,\texttt{<STOP>}$
@@
$\texttt{<think>}\,E\,\texttt{</think>}\quad\texttt{<answer>}\,y\,\texttt{</answer>}$
@@
where $x_{i}$ denotes the $i$-th token in the assistant response. Only the tokens in $X_{\text{a}}$ are used to compute the loss. This tuning step encourages the model to align with the structure and reasoning required for misinformation detection.
Stage 3: GRPO-based Reasoning Optimization. To enhance reasoning and explanation quality for multimodal misinformation detection, we adopt Group Relative Policy Optimization (GRPO) [shao2024deepseekmath]. GRPO ranks candidate outputs within each batch, which fits our setting where explanations for subtle visual edits and bilingual subtitle inconsistencies may be partially correct, and avoids the reward-model sensitivity of PPO and the pairwise preference assumption of DPO.
@@
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta_{\text{old}}}(q)}\bigg[\frac{1}{N}\sum_{i=1}^{N}\Big(\min\!\big(s_{1}A_{i},\;s_{2}A_{i}\big)$ (2)
$\qquad-\beta\,\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\big)\Big)\bigg].$
$A_{i}$ $=\frac{r_{i}-\mathrm{mean}(r_{1},\ldots,r_{N})}{\mathrm{std}(r_{1},\ldots,r_{N})},$ (3)
$s_{1}$ $=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\qquad s_{2}=\mathrm{clip}\!\left(s_{1},\;1-\epsilon,\;1+\epsilon\right).$
@@
Format reward. To enforce structured outputs, we define a format reward $R_{\text{format}}$ that equals 1 when the model output follows the predefined format with <answer></answer> and optional <think></think> tags, and 0 otherwise: $R_{\text{format}}=\mathds{1}\!\left[\text{output matches expected format}\right].$ Localization reward. We define an IoU-based reward $R_{\text{Loc}}=\frac{|\mathcal{M}_{\text{pred}}\cap\mathcal{M}_{\text{gt}}|}{|\mathcal{M}_{\text{pred}}\cup\mathcal{M}_{\text{gt}}|}$, where $\mathcal{M}_{\text{pred}}$ and $\mathcal{M}_{\text{gt}}$ are the predicted and ground-truth bounding boxes of the tampered region. Classification reward. We define $R_{\text{cls}}=\mathds{1}\!\left[C_{\text{pred}}=C_{\text{gt}}\right].$ to give a reward of 1 when the predicted label matches the ground truth and 0 otherwise.
The final reward combines all task-specific objectives, including formatting, classification, localization: $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$. This unified reward encourages the model to generate structured, accurate, and interpretable predictions across modalities without introducing any additional weighting coefficients, so each objective contributes equally.
@@ 5.1 Experimental Setup
Table 3: Combined results on BiMiBench and MMFakeBench (metrics in %). IoU is reported when region-level localization is available (“–” denotes not applicable). All results are averaged over three runs with standard deviation below 0.5%. Bold and underline indicate best and second-best performance, respectively.
Method MLLM BiMiBench MMFakeBench
ACC F1 IoU BERT ACC F1
InternVL3 [zhu2025internvl3exploringadvancedtraining] Qwen2.5-7B 20.85 (-25.95) 18.42 (-24.37) 7.23 (-23.01) 71.82 (-11.08) 42.69 (-28.25) 34.80 (-27.69)
InternVL3 [zhu2025internvl3exploringadvancedtraining] Qwen2.5-14B 28.47 (-18.33) 26.71 (-16.08) 14.38 (-15.86) 76.54 (-6.36)
Qwen3 [qwen3] Qwen3-8B 21.43 (-25.37) 14.96 (-27.83) 6.28 (-23.96) 73.92 (-8.98) 44.82 (-26.12) 39.49 (-23.00)
LLaVA-1.6 [liu2024llavanext] Vicuna-7B 22.91 (-23.89) 20.74 (-22.05) 8.91 (-21.33) 75.77 (-7.13) 30.09 (-40.85) 24.31 (-38.18)
Llama 3.1 [grattafiori2024llama3herdmodels] Llama-3.1-8B 24.61 (-22.19) 21.94 (-20.85) 5.27 (-25.0) 72.91 (-9.99) 32.67 (-38.27) 28.30 (-34.19)
Gemma3 [team2025gemma] Gemma3-4B-IT 17.48 (-29.32) 11.73 (-31.06) 4.29 (-25.95) 70.84 (-12.06) 27.19 (-43.75) 22.04 (-40.45)
SNIFFER [qi2024sniffer] Vicuna-13B 37.90 (-8.90) 33.63 (-9.16) 80.41 (-2.49) 67.49 (-3.45) 61.44 (-1.05)
MMD-Agent [liu2024mmfakebench] Qwen2.5-7B 34.78 (-12.02) 29.48 (-13.31) 62.78 (-8.16) 52.83 (-9.66)
BiMi (Ours) Gemma3-4B-IT 46.80 42.79 30.24 82.90 70.94 62.49
@@
We evaluate BiMi on BiMiBench and MMFakeBench [liu2024mmfakebench]. On BiMiBench, the model performs six-class misinformation classification, region-level tamper localization, and explanation generation. For MMFakeBench, which does not provide subtitles, we follow its standard four-class setting and supply minimal textual context via a unified prompt; subtitle-specific objectives are disabled while the rest of the pipeline remains unchanged.
Implementation Details. We adopt Gemma-3 [team2025gemma] as the base multimodal large language model and apply the proposed three-stage post-training pipeline: domain align, supervised fine-tuning (SFT) and GRPO-based reinforcement learning. The vision encoder is kept frozen throughout post-training, while only the multimodal adapter and language-model parameters are updated. All experiments are conducted on 48 GB NVIDIA V100 GPUs with a global batch size of 16 and an initial learning rate of $1\times 10^{-5}$. SFT is run for 3 epochs with early stopping based on validation loss, followed by 2 additional epochs of GRPO training. We fix the random seed to 42 and use AdamW with a weight decay of $1\times 10^{-2}$. The final reward combines the format, classification, and localization components without introducing any weighting coefficients, i.e., $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$, so that each objective contributes equally.
@@
Baselines. We compare BiMi against a range of strong multimodal baselines, including InternVL3 [zhu2025internvl3exploringadvancedtraining], Qwen3 [qwen3], LLaVA-1.6 [liu2024llavanext], and LLama3.1 [grattafiori2024llama3herdmodels], which represent leading approaches in vision-language modeling. We also include misinformation detection systems such as SNIFFER [qi2024sniffer] and MMD-Agent [liu2024mmfakebench], designed for out-of-context and multimodal misinformation. All models are evaluated using the same inputs: the image with overlaid Chinese-English subtitles and a task-specific prompt.
Evaluation Metrics. We report Accuracy (ACC) and F1 for classification and IoU for tampered-region localization. For explanation quality, we follow Liu et al. [liu2023llava] and compute BERTScore [bert-score] between model outputs and pseudo-references generated by GPT-4o [hurst2024gpt]. Pseudo-references are produced via structured prompting with label-grounded templates and manually verified by five trained annotators to ensure factual correctness and strict alignment with the ground-truth manipulations. To confirm that BERTScore reflects human judgment, the same annotators also rated 400 randomly sampled explanations on a 5-point Likert scale across five ground-truth–aligned dimensions—subtitle alignment, visual detail consistency, reasoning consistency, clarity and readability, and completeness. The high agreement between these human ratings (Fleiss’ $\kappa=0.71$) demonstrates the reliability and practical relevance of our evaluation protocol. More evaluation details are provided in the Appendix.
@@ 5.2 Main Results
Performance Comparison. We evaluate BiMi across three key subtasks on BiMiBench—(1) misinformation detection, (2) tampered region localization, and (3) explanation generation—and additionally assess cross-dataset generalization on MMFakeBench. As summarized in Table 3, BiMi consistently outperforms all baseline MLLMs across every metric and benchmark. On BiMiBench, it achieves notable improvements over the strongest baseline, including a +8.9 gain in detection accuracy, a +15.9 improvement in localization IoU, and a +2.5 increase in BERTScore for explanation quality. These gains are further strengthened by our GRPO-based reward design, which enhances both localization precision and explanation specificity. BiMi also achieves the best performance on MMFakeBench (4-class classification task), demonstrating strong robustness and generalization beyond the training domain.
Computation and Efficiency. End-to-end latency includes three stages: model forward pass, OCR subtitle extraction, and external retrieval. Pure model inference averages  320 ms per sample, while enabling OCR and web search adds 0.8–1.2 s depending on network conditions—similar to typical retrieval-augmented MLLMs. To ensure robustness, BiMi adopts a graceful-degradation strategy: if OCR or retrieval times out, the system automatically falls back to internal reasoning with minimal performance impact.
Ablation Study. We investigate the contributions of each stage in our pipeline using an ablation study. (i) Domain Alignment. Removing this stage reduces accuracy by 2.35 and BERTScore by 4.48, indicating its role in grounding visual-textual reasoning on news content. (ii) Instruction Tuning. Eliminating instruction tuning leads to a drastic performance drop, accuracy falls by 31.18 and BERTScore by 12.06, highlighting its importance in enabling bilingual understanding. (iii) GRPO. Without GRPO, the model fails to generate specific explanations tied to image or
@@
subtitle content, reducing accuracy by 5.7 and BERTScore by 8.09. (iv) Output Order. Reversing the output order (classify before explain) slightly lowers accuracy (–1.94) and BERTScore (–1.11), suggesting that generating explanations facilitates more reasoning. Table 4 summarizes the results. These findings confirm that all components are essential, with instruction tuning and GRPO being particularly critical for boosting accuracy and explanation quality.
Figure 5: Comparison of explanations from InternVL3 (middle) and our model (right). Top left: input; bottom left: original sample. Some responses are truncated due to space constraints.
@@ 5.3 Explainability Analysis
Beyond accurate classification and localization, generating high-quality natural language explanations is crucial for interpretable misinformation detection. We analyze explanation ability through qualitative comparisons, the influence of the retrieval module, and evaluation under bilingual.
Explanation Quality. We compare BiMi with InternVL3 [zhu2025internvl3exploringadvancedtraining] on a representative case (Figure 5). Both models capture the bilingual subtitle alignment, but InternVL3 relies on external cues and misses the image–text mismatch. BiMi instead its decision in the image, correctly localizes the manipulation, and offers a cleare explanation.
Retrieval Module on Real Samples. We further tested the retrieval component on a 100-sample set of authentic bilingual news images to examine its behavior in real-world scenarios. Human evaluation shows that retrieved evidence improves explanation clarity or correctly resolves subtle cross-lingual inconsistencies in 9% of cases, remains neutral in 89%, and results in off-topic passages in only 2%, typically due to OCR noise in Chinese subtitles. Importantly, even in these rare failure cases, the final predictions remain stable, indicating that retrieval errors do not propagate into the core reasoning pipeline. Overall, these results suggest that the retrieval module provides timely factual context that can meaningfully enrich the model’s explanations, while its downside risk is minimal—an encouraging property for practical deployment in noisy real-world environments.
@@
Bilingual multimodal capability. To assess bilingual understanding, we also create a single-language test variant where the model receives only Chinese or only English subtitles, reducing the task to a 4-class classification. Removing cross-lingual comparison makes the task substantially easier. As shown in Table 5, performance is noticeably higher with English-only inputs, reflecting the dominance of English data and reasoning patterns in pre-training. Chinese-only inputs introduce greater variance, particularly with idioms or long, multi-clause subtitles. These results reveal a clear gap in bilingual robustness and underscore the need for stronger Chinese modeling, improved tokenization, and more balanced multilingual data in future systems.
Failure Cases. BiMi’s remaining errors stem mainly from two sources. First, extremely subtle manipulations can lead to off-target localization or generic explanations, especially when visual edits or cross-lingual inconsistencies are barely perceptible. Second, OCR quality has a substantial impact: about 37% of misclassifications result from truncated characters or rare-character errors in Chinese subtitles, issues rarely seen in English text. As shown in Table 6, performance drops noticeably when using OCR-extracted subtitles instead of ground-truth text, confirming OCR noise as a major contributor. More examples are in the Appendix.
Table 6: Classification accuracy (ACC) and BERTScore (%) using Ground-truth subtitles (manually provided) versus OCR-extracted subtitles (automatically recognized).
Language Subtitle Source ACC (%) BERTScore
English Ground-truth 84.24 88.10
English OCR-extracted 83.02 85.92
Chinese Ground-truth 80.07 86.54
Chinese OCR-extracted 75.33 80.21